# Jarník-Besicovitch type theorems for semisimple algebraic groups

### v1

@@ 3. Counting rational points
Let $U$ be a small open bounded subset in $R_{u}(\overline{\mathbf{P}}_{0})$. Then for any $j\in J$ with $x_{j}\in\mathcal{K}\cap\mathbf{G}(\mathbb{R})^{0}$ and any sufficiently large $l>0$, we have
$m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l}))\sim l^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U)$
and
$m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4}}(U,A_{l},B_{l}))\sim l^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U).$
@@
Remark 3.4.
For our purpose, in the following, we fix some compact subsets $K_{1}\subset\ker\beta_{0}$, $K_{2}\subset\mathbf{H}_{e}(\mathbb{R})$, $K_{3}\subset\mathbf{M}_{a}(\mathbb{R})\cap\mathbf{G}(\mathbb{R})^{0}$ and $K_{4}\subset R_{u}(\mathbf{P}_{0})(\mathbb{R})$ discussed above.
4. A lower bound for the Hausdorff dimension of $S_{\gamma}^{c}$
In this section, we prove Theorem 1.3 under the assumption that the action of $\{a_{t}\}_{t\in\mathbb{R}}$ on $\mathbf{G}(\mathbb{R})^{0}/\Gamma$ is mixing.
@@ 4. A lower bound for the Hausdorff dimension of $S_{\gamma}^{c}$
Fix $j\in J$ with $x_{j}\in\mathbf{G}(\mathbb{R})^{0}$. Let $U$ be an open bounded subset in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ and
@@
$a_{T}\cdot S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})\Gamma=a_{T}\cdot U\Gamma/\Gamma\cap\exp(\mathfrak{a}_{K_{1},I_{0}})K_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma.$
@@
$a_{T}\cdot B_{\mathbf{H}_{e}}(\theta_{1},T)\cdot p=B_{\mathbf{H}_{e}}(\theta_{1})\cdot a\cdot h\cdot m\cdot u\cdot\gamma\subset\exp(\mathfrak{a}_{I_{0},K_{1}})\tilde{K}_{2}K_{3}K_{4}x_{j}\Gamma$
@@
Lemma 4.2.
We have
$\max\{\alpha(a_{1}):\alpha\in\Phi(\mathbf{F}_{e})\}>0.$
Proof.
Suppose on the contrary that for all $\alpha\in\Phi(\mathbf{F}_{e})$, $\alpha(a_{1})=0$. Let $\Omega_{\mathbf{F}_{e}}$ and $\Omega_{\mathbf{H}_{e}}$ be small open neighborhoods of identity in $\mathbf{F}_{e}(\mathbb{R})$ and $\mathbf{H}_{e}(\mathbb{R})$ respectively, and
$\Omega_{R_{u}(\overline{\mathbf{P}}_{0})}:=\Omega_{\mathbf{F}_{e}}\cdot\Omega_{\mathbf{H}_{e}}.$
Note that $\mathbf{H}_{e}(\mathbb{R})$ stabilizes every element in $V_{\beta_{0}}$. Now for any $p=f\cdot h\in\Omega_{R_{u}(\overline{\mathbf{P}}_{0})}$ with $f\in\Omega_{\mathbf{F}_{e}}$ and $h\in\Omega_{\mathbf{H}_{e}}$, we have
$a_{t}\cdot p=(a_{t}fa_{-t})(a_{t}ha_{-t})a_{t}=f\cdot(a_{t}ha_{-t})\cdot a_{t}.$
The element $a_{t}$ is rational and
$\delta(\rho(a_{t})\cdot\mathbb{Z}^{d}\cap V_{\beta_{0}})=e^{\beta_{0}(a_{t})}.$
This implies that
$\delta(\rho(a_{t}\cdot p)\cdot\mathbb{Z}^{d})\leq\kappa\cdot e^{-\beta_{0}(a_{-1})t}$
for any $t>0$ where the constant $\kappa$ depends only on $\Omega_{\mathbf{F}_{e}}$. Let $\Omega_{\mathbf{P}_{0}}$ be a small neighborhood of identity in $\mathbf{P}_{0}$. Then one can deduce from the inequality above that for any point $p\Gamma\in\Omega_{\mathbf{P}_{0}}\cdot\Omega_{R_{u}(\overline{\mathbf{P}}_{0})}\Gamma$, the orbit $a_{t}\cdot p\Gamma$ diverges in $\mathbf{G}(\mathbb{R})^{0}/\Gamma$. This contradicts the mixing property of $\{a_{t}\}_{t\in\mathbb{R}}$ on $\mathbf{G}(\mathbb{R})^{0}/\Gamma$. ∎
Let $X$ be a Riemannian manifold, $m$ a volume form on $X$ and $E$ a compact subset of $X$. Denote by $\operatorname{diam}(S)$ the diameter of a set $S\subset X$. A collection $\mathcal{A}$ of compact subsets of $E$ is said to be tree-like if $\mathcal{A}$ is the union of finite sub-collections $\mathcal{A}_{k}$ such that
@@
    For any $k$ and $S_{1},S_{2}\in\mathcal{A}_{j}$, either $S_{1}=S_{2}$ or $S_{1}\cap S_{2}=\emptyset$;
@@
We write $\mathbf{A}_{k}:=\bigcup_{A\in\mathcal{A}_{k}}A$ and $\mathbf{A}_{\infty}:=\bigcap_{k\in\mathbb{N}}\mathbf{A}_{k}$. We also define
@@
Proof of Theorem 1.3.
We fix the compact subsets $K_{1}\subset\ker(\beta_{0}),K_{2}\subset\mathbf{H}_{e}(\mathbb{R}),K_{3}\subset\mathbf{M}_{a}(\mathbb{R})\cap\mathbf{G}(\mathbb{R})^{0},K_{4}\subset R_{u}(\mathbf{P}_{0})$ as in Remark 3.4, and also fix $j\in J$ with $x_{j}\in\mathcal{K}\cap\mathbf{G}(\mathbb{R})^{0}$. Let $\epsilon>0$ be a sufficiently small number. Let $\nu_{0}$ be the $\mathbb{Q}$-root in $R_{u}(\overline{\mathbf{P}}_{0})$ such that
@@
Then by Lemma 4.2, $\nu_{0}(a_{1})>0$.
We start with a small open bounded box $U_{0}$ in $R_{u}(\overline{\mathbf{P}}_{0})$.
@@
$m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}}))\sim l_{1}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U_{0})$
@@
$m_{\mathbf{H}_{e}}(S_{K_{1},\tilde{K}_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}}))\sim l_{1}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U_{0})$
where $\tilde{K}_{2}=B_{\mathbf{H}_{e}}(\theta_{2})\cdot K_{2}$ for some $\theta_{2}>0$ as defined in Lemma 4.1. Now for any leaf $\mathcal{F}_{q}=\mathbf{H}_{e}(\mathbb{R})\cdot q$ ($q\in\mathbf{F}_{e}(\mathbb{Q})$) in the foliation $\mathcal{F}_{\mathbf{H}_{e}}$ such that
@@
we devide the region $\mathcal{F}_{q}\cap U_{0}$ into small cubes of side length
@@
where $\theta_{1}>0$ is the constant defined in Lemma 4.1. Then we collect those cubes $R$ which intersect $S_{K_{1},K_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}})$, and denote the corresponding collection by
$\mathcal{G}_{1,q}=\{R:R\cap S_{K_{1},K_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}})\neq\emptyset\}.$
@@
where $\tilde{K}_{2}=B_{\mathbf{H}_{e}}(\theta_{2})\cdot K_{2}$. Let
$\mathcal{H}_{1}=\bigcup_{\mathcal{F}_{q}\cap S_{K_{1},K_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}})\neq\emptyset}\bigcup_{R\in\mathcal{G}_{1,q}}R.$
@@
$\sim m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}}))$
$\leq m_{\mathbf{H}_{e}}(\bigcup_{\mathcal{F}_{q}\cap S_{K_{1},K_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}})\neq\emptyset}\bigcup_{R\in\mathcal{G}_{1,q}}R)$
@@
$\sim l_{1}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U_{0})$
and
$m_{\mathbf{H}_{e}}(\mathcal{H}_{1})\sim l_{1}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U_{0}).$
@@
$B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\times\exp(\mathfrak{a}_{K_{1},I_{0}})\tilde{K}_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma\to B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\exp(\mathfrak{a}_{K_{1},I_{0}})\tilde{K}_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma$
@@
$a_{-t_{1}}\cdot B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\cdot a_{t_{1}}\subset B_{\mathbf{F}_{e}}(\tilde{\delta}_{0},T_{1}).$
@@
may be larger than the diameter of $U_{0}$, so some cubes we obtain here from dividing the subset $\mathcal{P}_{1}$ may be outside the set $U_{0}$. For our purpose, we collect only those cubes which are inside the subset $U_{0}$. In this manner, we obtain a family of disjoint cubes constructed from sets in $\mathcal{F}_{1}(U_{0})$ inside $U_{0}$, which we denote by $\mathcal{A}_{1}$.
We remark here that if $\Phi^{0}(\mathbf{F}_{e})=\emptyset$, then all the cubes we obtain from dividing $\mathcal{P}_{1}$ are inside the subset $U_{0}$ if $T_{1}>0$ and $l_{1}>0$ are chosen to be sufficiently large (and also if we shrink or enlarge $U_{0}$ slightly to avoid the complexity caused by the boundary of $U_{0}$). We will see later that in the computations there is little difference between the case $\Phi^{0}(\mathbf{F}_{e})\neq\emptyset$ and the case $\Phi^{0}(\mathbf{F}_{e})=\emptyset$. Indeed, when we apply Theorem 4.3 and calculate $\Delta_{i}(\mathcal{A})$ and $d_{i}(\mathcal{A})$, the difference between these two cases may affect the value of the formula in finite steps, but eventually when we take the limit, this difference will disappear since we keep choosing sufficiently large numbers $T_{1},l_{1}$ and later $T_{2},l_{2},T_{3},l_{3},\dots$ to offset the effects by the difference at early stages.
Similarly, we can construct $\mathcal{A}_{k}$ inductively for any $k\in\mathbb{N}$. For $k>1$ we choose sufficiently large numbers $l_{k}>0$ and $T_{k}>0$ such that $\beta_{0}(a_{\tau_{k}})=-\ln l_{k}$. For each cube $S\in\mathcal{A}_{k-1}$, by the proposition above, we know that
$m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}}))\sim l_{k}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(S)$
@@
$m_{\mathbf{H}_{e}}(S_{K_{1},\tilde{K}_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}}))\sim l_{k}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(S)$
@@
we devide the region $\mathcal{F}_{q}\cap S$ into small cubes of side length
@@
where $\theta_{1}>0$ is the constant defined in Lemma 4.1. Then we collect those cubes which intersect $S_{K_{1},K_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}})$, and denote the corresponding collection by
$\mathcal{G}_{k,q,S}=\{R:R\cap S_{K_{1},K_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}})\neq\emptyset\}.$
@@
where $\tilde{K}_{2}=B_{\mathbf{H}_{e}}(\theta_{2})\cdot K_{2}$. Let
$\mathcal{H}_{k,S}=\bigcup_{\mathcal{F}_{q}\cap S_{K_{1},K_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}})\neq\emptyset}\bigcup_{R\in\mathcal{G}_{k,q,S}}R.$
@@
$\sim m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}}))$
$\leq m_{\mathbf{H}_{e}}(\bigcup_{\mathcal{F}_{q}\cap S_{K_{1},K_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}})\neq\emptyset}\bigcup_{R\in\mathcal{G}_{k,q,S}}R)$
@@
$\sim l_{k}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(S)$
and
$m_{\mathbf{H}_{e}}(\mathcal{H}_{k,S})\sim l_{k}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(S).$
@@
$B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\times\exp(\mathfrak{a}_{K_{1},I_{0}})\tilde{K}_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma\to B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\exp(\mathfrak{a}_{K_{1},I_{0}})\tilde{K}_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma$

### v2

@@ 3. Counting rational elements
Let $U$ be a small open bounded subset in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ which projects injectively into $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})/R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})\cap\Gamma$. Then for any $j\in J$ with $x_{j}\in\mathcal{K}\cap\mathbf{G}(\mathbb{R})^{0}$, we have
$m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l}))\asymp l^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U)$
as $l\to\infty$. Consequently, we have
$m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4}}(U,A_{l},B_{l}))\asymp l^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U).$
@@
4. Proof of Theorem 1.4: The case where $0\leq\tau(\psi)<\beta_{0}(a_{-1})$ and the flow $\{a_{t}\}_{t\in\mathbb{R}}$ is mixing
In this section, we prove Theorem 1.4 under the assumption that $0\leq\tau(\psi)<\beta_{0}(a_{-1})$ and the action of $\{a_{t}\}_{t\in\mathbb{R}}$ on $\mathbf{G}(\mathbb{R})^{0}/\Gamma$ is mixing. In the following, we fix some compact subsets $K_{1}\subset\ker\beta_{0}$, $K_{2}\subset\mathbf{H}_{e}(\mathbb{R})$, $K_{3}\subset\mathbf{M}_{a}(\mathbb{R})\cap\mathbf{G}(\mathbb{R})^{0}$ and $K_{4}\subset R_{u}(\mathbf{P}_{0})(\mathbb{R})$.
@@
Fix $j\in J$ with $x_{j}\in\mathbf{G}(\mathbb{R})^{0}$. Let $U$ be a small open bounded subset in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ and
@@
$a_{T}\cdot B_{\mathbf{H}_{e}}(\theta_{1},T)\cdot p=B_{\mathbf{H}_{e}}(\theta_{1})\cdot a\cdot h\cdot m\cdot u\cdot x_{j}\cdot\gamma\subset\exp(\mathfrak{a}_{I_{0},K_{1}})\tilde{K}_{2}K_{3}K_{4}x_{j}\Gamma$
@@
Let $X$ be a Riemannian manifold, $m$ a volume form on $X$ and $E$ a compact subset of $X$. Denote by $\operatorname{diam}(S)$ the diameter of a set $S\subset X$. A collection $\mathcal{A}$ of compact subsets of $E$ is said to be tree-like if $\mathcal{A}$ is the union of finite sub-collections $\mathcal{A}_{k}$ $(k\in\mathbb{N})$ such that
@@
    For any $k$ and $S_{1},S_{2}\in\mathcal{A}_{k}$, either $S_{1}=S_{2}$ or $S_{1}\cap S_{2}=\emptyset$;
@@
We write $\mathbf{A}_{k}:=\bigcup_{A\in\mathcal{A}_{k}}A$ and define a Cantor-type subset $\mathbf{A}_{\infty}:=\bigcap_{k\in\mathbb{N}}\mathbf{A}_{k}$. We also define
@@
We will use the following result to compute the lower bound of $\dim_{H}S_{\rho}(\psi)^{c}$ in Theorem 1.4.
@@
Proof of Theorem 1.4 assuming that $0\leq\tau(\psi)<\beta_{0}(a_{-1})$ and the flow $\{a_{t}\}_{t\in\mathbb{R}}$ is mixing.
We first get a lower bound for the Hausdorff dimension of $S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. We fix some compact subsets $K_{1}\subset\ker(\beta_{0}),K_{2}\subset\mathbf{H}_{e}(\mathbb{R}),K_{3}\subset\mathbf{M}_{a}(\mathbb{R})\cap\mathbf{G}(\mathbb{R})^{0}$ and $K_{4}\subset R_{u}(\mathbf{P}_{0})(\mathbb{R})$ as in the begining of this section, and also fix some index $j\in J$ with $x_{j}\in\mathcal{K}\cap\mathbf{G}(\mathbb{R})^{0}$. Assume that $0\leq\tau=\tau(\psi)<\beta_{0}(a_{-1})$. Let $\epsilon>0$ be a sufficiently small number such that $\tau+\epsilon<\beta_{0}(a_{-1})$. Let $\nu_{0}$ be any $\mathbb{Q}$-root relative to $\mathbf{T}$ in $R_{u}(\overline{\mathbf{P}}_{0})$ such that
@@
Since the unstable horospherical subgroup of $\{a_{t}\}_{t\in\mathbb{R}}$ is contained in $R_{u}(\overline{\mathbf{P}}_{0})$, we have $\nu_{0}(a_{1})>0$.
In the following, we construct a tree-like collection $\mathcal{A}=\{\mathcal{A}_{k}\}_{k\in\mathbb{N}}$ of compact subsets in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. We start with a small open bounded box $U_{0}$ in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ which projects injectively into $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})/(R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})\cap\Gamma)$.
@@
$m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}}))\asymp l_{1}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U_{0})$
@@
$m_{\mathbf{H}_{e}}(S_{K_{1},\tilde{K}_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}}))\asymp l_{1}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U_{0})$
where $\tilde{K}_{2}=B_{\mathbf{H}_{e}}(\theta_{2})\cdot K_{2}$ for some $\theta_{2}>0$ defined in Lemma 4.1. Now for any leaf $\mathcal{F}_{q}=\mathbf{H}_{e}(\mathbb{R})\cdot q$ ($q\in\mathbf{F}_{e}(\mathbb{Q})$) in the foliation $\mathcal{F}_{\mathbf{H}_{e}}$ such that
@@
we divide the region $\mathcal{F}_{q}\cap U_{0}$ into small cubes of side length
@@
where $\theta_{1}>0$ is the constant defined in Lemma 4.1. Then we collect those cubes $R\subset\mathcal{F}_{q}\cap U_{0}$ which intersect $S_{K_{1},K_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}})$, and denote the corresponding collection by
$\mathcal{G}_{1,q}:=\{R:R\subset\mathcal{F}_{q}\cap U_{0},\;R\cap S_{K_{1},K_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}})\neq\emptyset\}.$
@@
where $\tilde{K}_{2}=B_{\mathbf{H}_{e}}(\theta_{2})\cdot K_{2}$. For any leaf $\mathcal{F}_{q}=\mathbf{H}_{e}(\mathbb{R})\cdot q$ $(q\in\mathbf{F}_{e}(\mathbb{Q}))$ in $\mathcal{F}_{\mathbf{H}_{e}}$ such that
$\mathcal{F}_{q}\cap S_{K_{1},K_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}})=\emptyset,$
we set $\mathcal{G}_{1,q}=\emptyset$.
Now let
$\mathcal{H}_{1}=\bigcup_{q\in\mathbf{F}_{e}(\mathbb{Q})}\bigcup_{R\in\mathcal{G}_{1,q}}R.$
@@
$\asymp m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(U_{0},A_{l_{1}},B_{l_{1}}))$
$\leq m_{\mathbf{H}_{e}}(\bigcup_{q\in\mathbf{F}_{e}(\mathbb{Q})}\bigcup_{R\in\mathcal{G}_{1,q}}R)$
@@
$\asymp l_{1}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U_{0})$
and hence
$m_{\mathbf{H}_{e}}(\mathcal{H}_{1})\asymp l_{1}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U_{0}).$
@@
$B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\times\exp(\mathfrak{a}_{I_{0},K_{1}})\tilde{K}_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma\to B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\exp(\mathfrak{a}_{I_{0},K_{1}})\tilde{K}_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma$
@@
$a_{-t_{1}}\cdot B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\cdot a_{t_{1}}\subset B_{\mathbf{F}_{e}}(\tilde{\delta}_{0},T_{1})\textup{ and }a_{T_{1}}\cdot q\Gamma\in\exp(\mathfrak{a}_{K_{1},I_{0}})\tilde{K}_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma\;(\forall q\in\mathcal{H}_{1}).$
@@
may be larger than the diameter of $U_{0}$, so some cubes we obtain here (by dividing the subset $\mathcal{P}_{1}$) may be outside the set $U_{0}$. For our purpose, we collect only those cubes which are inside the subset $U_{0}$. In this manner, we obtain a family of disjoint cubes inside $U_{0}$ constructed from $\mathcal{P}_{1}$, which we denote by $\mathcal{A}_{1}$.
We remark here that if $\Phi^{0}(\mathbf{F}_{e})=\emptyset$, then all the cubes we obtain by dividing $\mathcal{P}_{1}$ are inside the subset $U_{0}$ if $T_{1}>0$ and $l_{1}>0$ are chosen to be sufficiently large (and also if we enlarge $U_{0}$ slightly to avoid the complexity caused by the boundary of $U_{0}$). We will see later that in the computations there are no significant differences between the case $\Phi^{0}(\mathbf{F}_{e})\neq\emptyset$ and the case $\Phi^{0}(\mathbf{F}_{e})=\emptyset$. Indeed, when we apply Theorem 4.2 and calculate $\Delta_{i}(\mathcal{A})$ and $d_{i}(\mathcal{A})$, the differences between these two cases may affect the value of the formula in finite steps, but eventually when we take the limit, these differences will disappear since we keep choosing sufficiently large numbers $T_{1},l_{1},T_{2},l_{2},T_{3},l_{3},\dots$ to offset the effects by the differences in the early stages.
In a similar way, we can construct $\mathcal{A}_{k}$ for any $k\in\mathbb{N}$ inductively. For $k>1$ we choose sufficiently large numbers $l_{k}>0$ and $T_{k}>0$ such that $\beta_{0}(a_{T_{k}})=-\ln l_{k}$. For each cube $S\in\mathcal{A}_{k-1}$, by Proposition 3.3, we know that
$m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}}))\asymp l_{k}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(S)$
@@
$m_{\mathbf{H}_{e}}(S_{K_{1},\tilde{K}_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}}))\asymp l_{k}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(S)$
@@
we divide the region $\mathcal{F}_{q}\cap S$ into small cubes of side length
@@
where $\theta_{1}>0$ is the constant defined in Lemma 4.1. Then we collect those cubes $R\subset\mathcal{F}_{q}\cap S$ which intersect $S_{K_{1},K_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}})$, and denote the corresponding collection by
$\mathcal{G}_{k,q,S}=\{R:R\subset\mathcal{F}_{q}\cap S,\;R\cap S_{K_{1},K_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}})\neq\emptyset\}.$
@@
where $\tilde{K}_{2}=B_{\mathbf{H}_{e}}(\theta_{2})\cdot K_{2}$. For any leaf $\mathcal{F}_{q}=\mathbf{H}_{e}(\mathbb{R})\cdot q$ $(q\in\mathbf{F}_{e}(\mathbb{Q}))$ in $\mathcal{F}_{\mathbf{H}_{e}}$ such that
$\mathcal{F}_{q}\cap S_{K_{1},K_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}})=\emptyset,$
we set $\mathcal{G}_{k,q,S}=\emptyset$.
Let
$\mathcal{H}_{k,S}=\bigcup_{q\in\mathbf{F}_{e}(\mathbb{Q})}\bigcup_{R\in\mathcal{G}_{k,q,S}}R.$
@@
$\asymp m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(S,A_{l_{k}},B_{l_{k}}))$
$\leq m_{\mathbf{H}_{e}}(\bigcup_{q\in\mathbf{F}_{e}(\mathbb{Q})}\bigcup_{R\in\mathcal{G}_{k,q,S}}R)$
@@
$\asymp l_{k}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(S)$
and hence
$m_{\mathbf{H}_{e}}(\mathcal{H}_{k,S})\asymp l_{k}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(S).$
@@
$B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\times\exp(\mathfrak{a}_{I_{0},K_{1}})\tilde{K}_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma\to B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\exp(\mathfrak{a}_{I_{0},K_{1}})\tilde{K}_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma$